# Shaping Sparse Rewards in Reinforcement Learning: A Semi-supervised Approach

### v1

@@ C Overall Framework
$Q(s,a)\leftarrow Q(s,a)+\alpha\cdot\frac{1}{|\mathcal{B}|}\sum_{(s,a,\hat{r},s^{\prime})\in\mathcal{B}}\left(\hat{r}+\gamma\max_{a}Q(s^{\prime},a)-Q(s,a)\right).$
14:     until $s_{t+1}$ is the terminal state
15:     Update parameter $\theta$ through minimizing the objective $L=L_{QV}+\alpha L_{s}+(1-\alpha)L_{r}$
16:end for
Figure 1: SSRS framework. $r_{t}=0$ indicating the sparse reward samples in the buffer and $r_{t}\neq 0$ indicating samples with non-zero rewards revise the figure a bit.
Before updating the policy gradient or the Q-network parameters in the deep RL algorithm, the reward estimation network is used to sample sparse reward trajectories from the experience replay buffer with probability $p_{u}$ to compute the estimated reward. In the early stages of training, the update probability $p_{u}$ is set proportional to $\log(\mu\mathcal{N})$, where $\mu\mathcal{N}$ represents the number of trajectories containing true reward signals. This prevents excessive variance in the initial policy gradient updates or Q-network updates, which is crucial for maintaining stability during the training process. In the middle stages of training, the update probability $p_{u}$ is increased to match $\mu\mathcal{N}$, and in the later stages, it is reduced again to be proportional to $\log(\mu\mathcal{N})$, as the model has already learned a good policy and no longer requires a large number of simulated reward signals to assist in descent.
In addition, in the early stages of training, the model relies more on the true rewards from the environment feedback. As training progresses, the model will output reward estimates with increasingly higher confidence, i.e., satisfying $\max(\mathbf{q_{t}})\geq\lambda$, while also learning more domain knowledge of the trajectory space with increasing coefficient $\alpha$.
4 Experiment Results
The experiments aim to validate the performance of the SSRS framework in both the Atari game environment Bellemare u. a. (2013) and the robotics manipulation environment Plappert u. a. (2018a) (see Figure 2) and, using Random Access Memor (RAM) observations as an example, demonstrate the superiority of double entropy data augmentation over other data augmentation methods. Additionally, ablation experiments are conducted to demonstrate the importance of monotonicity constraints in the application of SSL techniques. Distinct from other reward shaping methods, our paper introduces an semi-supervised pipeline to enhance sample efficiency. In Section 4.4, we will analyze the characteristics of the trajectory space through experiments to provide insights into the underlying reasons of such improvement.
@@ 4 Experiment Results
Figure 2: (a) and (b) show the game scenes from Montezuma Revenge and Pitfall in Sparse reward environment of the Arcade Learning Environment (ALE), commonly referred to as Atari. (c) is a snapshot of robotic arm performing Fetch and Reach task in Gymnasium-Robotics environment.
Experimental Setup. RAM observation refers to a representation of the Atari RL environments’ internal state directly from its Random Access Memory, typically 128 bytes of RAM (i.e., a vector of 128 integers, each in the range [0, 255]). The code for this experiment is based on the Tianshou RL library Weng u. a. (2022). At each epoch, 2000 transition samples are collected (i.e., 2000 timesteps are executed in the environment) and appended to the dataset $\mathcal{D}$, a buffer with a capacity of 30k transitions. Some of the hyperparameter settings are as Table 3 in Appendix, which are determined based on preliminary hyperparameter tuning results, and consistent hyperparameter settings are used in both the Robotics and Atari experiments.
4.1 Performance of SSRS
Figure 3: this consensus matrix of reward distribution in Hero-ram-v4 after 1000 episodes iterations indicates that the distribution of trajectories along the reward dimension in the regulatory space exhibits a certain clustering property. The cluster method used is Gauss Mixture Models, and the number of iterations for consesus matrix is 100.
To evaluate the proposed SSRS framework, we compare it with several baselines and benchmarks algorithms in both the Atari and Robotic environments:
  • Intrinsic Curiosity Module (ICM, Pathak u. a. (2017)): ICM encourages exploration by rewarding the agent for discovering novel states based on intrinsic motivation.
  • Reward Conditioned Policy (RCP, Kumar u. a. (2019)): RCP conditions views non-expert trajectories collected from suboptimal policies as optimal supervision so as to shape the reward of the given trajectory.
  • Prioritized Experience Replay (PER, Schaul u. a. (2015)): PER improves learning efficiency by prioritizing experience samples with high temporal difference errors.
The baseline algorithms for both Atari and Robotic environments are Deep Q-Network (DQN, Mnih u. a. (2015)) and Deep Deterministic Policy Gradient (DDPG, Lillicrap u. a. (2015)), respectively. Several variants of the SSRS framework are employed, including SSRS-S, which uses double entropy augmentation, SSRS-C, which adopts the cutout method for data augmentation, and SSRS-M, which adopts the smooth method for data augmentation.
In the FetchReach environment in Robotics, where the reward signal is binary, i.e., the reward is -1 if the end effector hasn’t reached its final target position, and 0 if the end effector is in the final target position, the SSRS framework outperforms both PER and DDPG (see Figure 4), with the latter two algorithms struggling to learn a successful policy. However, the ICM algorithm performs better than SSRS in this environment by more effectively shaping the reward, achieving a higher best reward. In another sparse while more diverse environment, i.e., Atari (see Figure 5), the SSRS algorithm and its variants outperform the baseline algorithm across all four games. In games with moderate reward sparsity, such as Seaquest and Hero, SSRS achieves performance comparable to the ICM algorithm. In terms of convergence speed, the SSRS-S variants outperform ICM in the Seaquest environment. Moreover, in games with extremely sparse rewards, such as MonteZumaRevenge and Venture, SSRS significantly outperforms the ICM algorithm, achieving nearly 4 times the best score in MonteZumaRevenge (see Figure 5(c)). Evidently, the sparser the rewards in the environment, the more unstably the policy updates, as indicated by the larger variance in the best reward curves. In contrast, SSRS and its variants exhibit superior update stability in Venture during later stages, as shown in Figure 5(d).
Figure 4: The best score curve of SSRS variants and aforementioned baselines over 500 epochs in the robotic manipulation environment FetchReach.
4.2 Performance of Data Augmentation
(a) Seaquest (SSRS)
(b) Hero (SSRS)
@@ 4.2 Performance of Data Augmentation
Figure 5: Figures (a–d) demonstrate the best score in 1000 epochs of different algorithms in 4 Atari environments. Different curves represent the SSRS framework and its variants, as well as ICM, RCP and DQN algorithms. The legends of each sub figure are uniformly labeled in Figure(a)
(a) Seaquest (DA)
(b) Hero (DA)
(c) Montezuma’s Revenge (DA)
(d) Venture (DA)
Figure 6: Figures (a–d) demonstrate the best score in 1000 epochs of different data augmentation methods in 4 Atari environments. Note that Shannon refers to double entropy data augmentation. The legends of each sub figure are uniformly labeled in Figure(a)
The performance analysis of double entropy data augmentation is conducted in Atari environment (see Figure 6). In the first two games, Seaquest and Hero, double entropy data augmentation consistently maintains a leading position in terms of both convergence speed and final highest rewards, with the variance of the best reward remaining within a narrow range. In the more reward-sparse environment of MonteZumaRevenge, although the final best reward achieved by the translate augmentation (blue curve) method is higher, it exhibits significantly larger variance. In contrast, double entropy data augmentation (green curve) demonstrates much smaller variance (see Figure 6(c)), indicating a more stable policy and a faster update rate. As we will mention in Section 4.4, double entropy data augmentation preserves the smoothness and clustering of trajectories better in the trajectory space. This is crucial to the step of calculating estimated reward, where the $V$ network and $Q$ network are not able to separate trajectories to learn a good representation of the space. From the perspective of the teacher-student model, such a ‘muddled’ teacher would lead the student, i.e. policy module, to learn an unstable policy.
In Venture, double entropy data augmentation maintains stable variance throughout (see Figure 6(d)), whereas other data augmentation methods struggle to avoid larger variance in such a sparse environment, where disturbance to the data exceeded the decision-making boundaries of the trajectory space. Moreover, double entropy data augmentation achieves a best score that was 10 percentage points higher than the best result among tested image-based data augmentation methods in this environment.
4.3 Ablation Experiments of Monotonicity Constraint
@@
The average best score obtained is presented in Table 1. In all the three environments, SSRS with a monotonicity constraint achieves higher best scores compared to SSRS without a monotonicity constraint. Considering the results from previous two sections, as SSRS-S set (green curves) only differs in monotonicity constraints between SSRS experiments (see Figures 5) and DA experiments (see Figures 6), we find that adding the monotonicity constraint reduces the variance as well as the final best score in reward-sparse environments with long-term goals (see Figure 5(d) and 6(d)). However in the less reward-sparse Seaquest environment, this constraint increases the variance and the final best score (see Figure 5(a) and 6(a)). The lower variance means a more stable update in policy, but may also reflect an insufficient exploration, resulting in a lower best score compared to the unconstrained case, as observed in Table 1 for the Venture environment. One of the reasons of such difference in sparse and less-sparse scenarios is our strategy of configure in update probability $p_{u}$. The hyperparameter $p_{u}$ is set proportional to the number of non-zero-reward transitions, $\mu\mathcal{N}$. However, the variation in the number of non-zero-reward transitions across different environments is not linear. Although we improved the method by setting $p_{u}$ to the order of $\log(\mu\mathcal{N})$ during the early stages of training, this adjustment does not cover all possible scenarios.
Table 1: Average best score of SSRS framework with monotonicity (SSRS-ST) constraint and without monotonicity constraint (SSRS-NST) in 4 environments.
avg best score SSRS-ST SSLRL-NST
Seaquest 480 447.5
Hero 8870 7599.5
Montezuma 166.3 100
Venture 200 275
@@
Further investigation of the reward distribution in Hero (see Figure 7) reveals that, according to the reward distribution of the baseline algorithm without reward shaping, most of the reward signals in this environment consist of small reward values, with a lack of medium to long-term reward feedback. SSRS without a monotonicity constraint only provides reward estimates close to the mean. This may be related to the observation in Section 4.4 that reward values near the mean exhibit better clustering properties. However, this completely deviates from the actual reward distribution of the environment, causing the reward estimator to degrade into a binary classifier. Experimental results also indicate a performance drop without the constraint. In contrast, SSRS with a monotonicity constraint shapes a reward distribution that closely approximates the true reward distribution while also exhibiting the properties of a positively skewed normal distribution.
4.4 Feasibility of Semi-Supervised Learning
@@
SSL methods relies on the smoothness assumption and clustering assumption Ouali u. a. (2020) on the augmented data, in this case, the trajectories. To be specific, the smoothness assumption indicates if two points $x_{1},x_{2}$ reside in a high-density region are close, then so should be their corresponding outputs $y_{1},y_{2}$. And clustering assumption indicates that samples from the same class are closer to each other, while there are significant boundaries between samples from different classes and the decision boundary of the model should be as far as possible from regions of high data density. We examine the trajectory distribution and obtain the consensus matrix $\mathcal{C}$ in Figure 3. Each element $a_{ij}$ $\in\mathcal{C},0\leq a_{ij}\leq 1,$ is frequency at which the two trajectories $\tau_{i}$ and $\tau_{j}$ are assigned to the same shaped reward value. Note that we use reward shaped values for labels.
@@
Larger reward values, i.e., rewards given for achieving long-term goals, exhibit clear decision boundaries with smaller reward values. However, the decision boundaries inside high reward values are more ambiguous, with only one clear diagonal block, aligning with the clustering assumption. But this ambiguity has minimal impact on SSRS, as the the probabilities of reward estimator output these rewards with high confidence is negligible. By also referring to Figure 7, it can be observed that the three ‘phases’ of significance in the consensus matrix plot correspond to the downward trend of the reward distribution probability in Figure 7. This can be understood from the perspective that as the reward value increases, the randomness of the trajectory decreases, leading to greater clustering significance in the trajectory space, and vice versa.
5 Conclusion
@@
In this paper, we propose the Semi-Supervised Reward Shaping (SSRS) framework which utilizes zero reward trajectories by employing SSL technique. Additionally, we introduce the double entropy data augmentation method for consistency regularization and apply a monotonicity constraint over modules of reward estimator. Our model outperforms RCP and achieves performance comparable to ICM in Atari environments, with a maximum of fourfold increase in reaching higher best scores. Moreover, the double entropy data augmentation enhances performance showcasing a 15.8% increase in best score compared to other augmentation methods. With SSL techniques, we can deploy agents on significantly sparser trajectory data. From a reverse perspective, it is possible to obtain learnable trajectories by artificially annotating rewards for a minimal number of transitions within a large set of trajectory data.
@@
However, it’s important to note that SSRS also introduces several hyperparameters, such as the update probability of trajectories in the experience replay buffer $p_{u}$, which has a significant impact on algorithm performance. The tuning of $p_{u}$ reflects the balance between exploration and exploitation in this class of reward shaping algorithms. Similar issues were also highlighted in Peng u. a. (2019); Kumar u. a. (2019), where the problem of controlling the reward shaping ratio arises when the reward shaping estimator is suboptimal. Like our paper, the balance shifts towards deciding whether the agent should exploit the reward from the estimator or prioritize exploration. We believe that a theoretical analysis of the dynamic relationship between the optimality of the estimator and the exploration-exploitation ratio would be a valuable direction for future work.
Appendix A The Gradients of Loss Functions
A.1 The Gradient of $L_{QV}$
For $L_{QV}$ mentioned in Section 3.2 in the following explicit form:

### v2

@@ Introduction
$Q(s,a)\leftarrow Q(s,a)+\alpha\cdot\frac{1}{|\mathcal{B}|}\\ \cdot\sum_{(s,a,\hat{r},s^{\prime})\in\mathcal{B}}\left(\hat{r}+\gamma\max_{a}Q(s^{\prime},a)-Q(s,a)\right).$
13:  until $s_{t+1}$ is the terminal state
14:  Update parameter $\theta$ through minimizing the objective $L=L_{QV}+\alpha L_{s}+(1-\alpha)L_{r}$
15:end for
Experiment Results
@@
Figure 2: (a) and (b) show the game scenes from Montezuma Revenge and Pitfall as Sparse reward environment of the Arcade Learning Environment (ALE), commonly referred to as Atari. (c) is a snapshot of robotic arm performing Fetch and Reach task in Gymnasium-Robotics environment.
(a) Seaquest
(b) Hero
@@
Figure 3: Figures (a–d) demonstrate the best score in 1000 epochs of different algorithms in 4 Atari environments (mean $\pm$ std over 3 seeds). Different curves represent the SSRS framework and its variants, as well as RCP and SAC algorithms. The legends of each sub figure are uniformly labeled in Figure(a)
The experiments aim to validate whether the SSRS using semi-supervised learning is more efficient in reward shaping than the self-supervised method, thereby improving the best score achieved by the agent in both the Atari game environment (Bellemare et al. 2013) and the robotics manipulation environment (Plappert et al. 2018) (see Figure 2). To demonstrate the superiority of double entropy data augmentation over other data augmentation methods in non-visual observation tasks, we use Random Access Memory (RAM) as the observation. Additionally, ablation experiments are conducted to demonstrate the importance of monotonicity constraints in the application of SSL techniques. We will analyze the characteristics of the trajectory space through experiments to provide insights into the underlying reasons of such improvement.
Experimental Setup. RAM observation refers to a representation of the Atari RL environments’ internal state directly from its Random Access Memory, typically 128 bytes of RAM (i.e., a vector of 128 integers, each in the range [0, 255]). The implementation for this experiment is based on the Tianshou RL library (Weng et al. 2022). Across three experimental seeds, a maximum of 2000 transition samples per episode are collected and appended to the buffer $\mathcal{D}$, a buffer with a capacity of 30k transitions. The maximum test score achieved from the start of training up to the current epoch, along with its variance across different experimental seeds, is recorded and plotted in the figures throughout this section. Some of the hyperparameter settings are as Table 4 in Appendix, which are determined based on preliminary hyperparameter tuning results, and consistent hyperparameter settings are used in both the Robotics and Atari experiments.
Performance of SSRS
Figure 4: This figure shows the best score curve of SSRS variants and aforementioned baseline over 500 epochs in the robotic manipulation environment FetchReach (mean $\pm$ std over 3 seeds).
To evaluate the proposed SSRS framework, we compare it with several baseline approaches including supervised-based methods (RCP (Kumar, Peng, and Levine 2019) and SORS (Memarian et al. 2021)) as well as backbone reinforcement learning algorithms in both Atari and robotic environments. Since this paper focuses on reward shaping, the backbone algorithms used are not the main emphasis. The backbone algorithms for Atari and Robotic environments are Soft Actor-Critic (SAC, Haarnoja et al. (2019)) and Deep Deterministic Policy Gradient (DDPG, Lillicrap et al. (2015)), respectively. Several variants of the SSRS framework are employed, including SSRS-S, which uses double entropy augmentation, SSRS-C, which adopts the cutout method for data augmentation, and SSRS-M, which adopts the smooth method for data augmentation.
In Atari environment, the SSRS algorithm and its variants outperform the baseline algorithm and supervised-based RCP across all four games (see Figure 3). This is because When the proportion of transitions with non-zero rewards is low (i.e., when $\mu$ is small), the self-supervised method has fewer samples containing ”ground truth labels” available for learning. In such cases, the SSRS approach, which leverages semi-supervised learning to infer trajectory decision boundaries, can estimate rewards from transitions more effectively than the self-supervised method. (The concept of trajectory decision boundaries will be discussed in section ’Feasibility of Semi-Supervised Learning’.) Seaquest and Hero games are considered to be moderately reward-sparse, so we add them to the setting to evaluate the general performance of SSRS. In terms of convergence speed, the SSRS-S variants outperform RCP in the Venture environment. Moreover, in games with extremely sparse rewards, such as MonteZumaRevenge, SSRS significantly outperforms the RCP algorithm, achieving nearly 2 times the best score in MonteZumaRevenge (see Figure 3(c)). In the FetchReach environment in Robotics, where the reward signal is binary, i.e., the reward is -1 if the end effector hasn’t reached its final target position, and 0 if the end effector is in the final target position, the SSRS framework outperforms the self-supervised method SORS (see Figure 4).
We also conducte experiments to evaluate the performance of SSRS when using different data augmentation methods as perturbations for SSL, aiming to investigate whether the proposed double entropy data augmentation provides greater improvement in terms of final best score for SSRS compared to other traditional data augmentation methods (e.g., flip, scale, translate). The detailed results and analysis are provided in Appendix.
Table 1: Average best score of SSRS framework with monotonicity (SSRS-ST) constraint and without monotonicity constraint (SSRS-NST) in 4 environments (mean $\pm$ std over 3 seeds).
avg best score SSRS-ST SSLRL-NST
Seaquest 476.6 $\pm$ 85 447.5 $\pm$ 18.9
Hero 8825 $\pm$ 2317 7596.6 $\pm$ 81.4
Montezuma 166.6 $\pm$ 104 100 $\pm$ 70.7
Venture 200 $\pm$ 0 262.5 $\pm$ 47.9
Ablation Experiments of Monotonicity Constraint
@@
The average best score obtained is presented in Table 1. In three environments, SSRS with a monotonicity constraint achieves higher best scores compared to SSRS without a monotonicity constraint. We find that adding the monotonicity constraint reduces the variance as well as the final best score in reward-sparse environments with long-term goals (see Figure 3 and 1 in Appendix). One of the reasons of such difference in sparse and less-sparse scenarios is our strategy of configure in update probability $p_{u}$. The hyperparameter $p_{u}$ is set proportional to the number of non-zero-reward transitions, $\mathcal{N}_{r}$. However, the variation in the number of non-zero-reward transitions across different environments is not linear. Although we improved the method by setting $p_{u}$ to the order of $\log(\mathcal{N}_{r})$ during the early stages of training, this adjustment does not cover all possible scenarios.
@@
Further investigation of the reward distribution in Hero (see Figure 5) reveals that, according to the reward distribution of the backbone algorithm without reward shaping, most of the reward signals in this environment consist of small reward values, with a lack of medium to long-term reward feedback. SSRS without a monotonicity constraint only provides reward estimates close to the mean. However, this deviates from the actual reward distribution of the environment, causing the reward estimator to degrade into a binary classifier. Experimental results also indicate a performance drop without the constraint. In contrast, SSRS with a monotonicity constraint shapes a reward distribution that closely approximates the true reward distribution while also exhibiting the properties of a positively skewed normal distribution.
Feasibility of Semi-Supervised Learning
Figure 6: This figure shows the consensus matrix of reward distribution in Hero after 1000 episodes iterations, which indicates that the distribution of trajectories along the reward dimension in the trajectory space exhibits a certain clustering property. The cluster method used is Gaussian Mixture Models, and the number of iterations for consesus matrix is 100.
To further investigate why SSL method can infer rewards from trajectories more effectively than self-supervised methods—and why SSL techniques can be applied to reward shaping, we conduct the following experiment and analyses. We examine the trajectory distribution and obtain the consensus matrix $\mathcal{C}$ in Figure 6. Each element $a_{ij}$ $\in\mathcal{C},0\leq a_{ij}\leq 1,$ is frequency at which the two trajectories $\tau_{i}$ and $\tau_{j}$ are assigned to the same shaped reward value. Note that we use estimated reward values for labels.
@@
Larger estimated reward values, i.e., rewards given for achieving long-term goals, exhibit clearer diagonal boundaries compared to smaller reward values. However, the decision boundaries inside high reward values are more ambiguous, with only one clear diagonal block, aligning with the clustering assumption of SSL. SSL methods relies on the smoothness assumption and clustering assumption (Yang et al. 2023) on the augmented data, in this case, the trajectories. To be specific, for classification models, the smoothness assumption indicates if two points $x_{1},x_{2}$ reside in a high-density region are close, then so should be their corresponding outputs $y_{1},y_{2}$. And clustering assumption indicates that samples from the same class are closer to each other, while there are significant boundaries between samples from different classes and the decision boundary of the model should be as far as possible from regions of high data density. Notably, this observed ambiguity in high reward values has minimal practical impact on SSRS, as the reward estimator outputs these high rewards with negligible confidence probabilities.
Unlike conventional data augmentation techniques that may introduce excessive perturbations to non-image data—violating the smoothness and clustering assumptions— double entropy augmentation better preserves these assumptions in trajectory space. This is crucial to the step of calculating estimated reward, where the $V$ network and $Q$ network are not able to separate trajectories to learn a good representation of the space. From the perspective of the teacher-student model (Tarvainen and Valpola 2017), such a ‘muddled’ teacher would lead the student, i.e. policy module, to learn an unstable policy.
@@
By also referring to Figure 5, it can be observed that the three ‘phases’ of significance in the consensus matrix plot correspond to the downward trend of the reward distribution probability. This can be understood from the perspective that as the estimated reward value increases, the randomness of the trajectory decreases, leading to greater clustering significance in the trajectory space, and vice versa.
Conclusion
@@
In this paper, we propose a Semi-Supervised Reward Shaping (SSRS) framework which utilizes zero reward trajectories by employing SSL technique. Additionally, we propose the double entropy data augmentation method for consistency regularization and a monotonicity constraint over modules of reward estimator. Our model outperforms previous supervised-based RCP and SORS methods in Atari environments, with a maximum of twice increase in reaching higher best scores. Moreover, the double entropy data augmentation enhances performance showcasing a 15.8% increase in best score compared to other augmentation methods. With SSL techniques, we can deploy agents on significantly sparser trajectory data. From a reverse perspective, it is possible to obtain learnable trajectories by artificially annotating rewards for a minimal number of transitions within a large set of trajectory data.
@@
Moreover, it’s worth noting that SSRS also introduces several hyperparameters, such as the update probability of trajectories in the experience replay buffer $p_{u}$. Although the tuning of these hyperparameters has been studied in the literature of SSL, these hyperparameters can still provide valuable insights for the field of reinforcement learning. In this class of reward shaping algorithms, the tuning of $p_{u}$ reflects the balance between exploration and exploitation in that, a higher $p_{u}$ encourages the agent to exploit knowledge learned from the reward function, while a lower $p_{u}$ prompts the agent to explore according to its original policy. Similar issues were also highlighted in Peng et al. (2019); Kumar, Peng, and Levine (2019), where the problem of controlling the reward shaping ratio arises. Like our paper, the balance shifts towards deciding whether the agent should exploit the reward from the estimator or prioritize exploration. We believe that a theoretical analysis of the dynamic relationship between the optimality of the estimator and the exploration-exploitation ratio would be a valuable direction for future work.
Appendix A Double Entropy Data Augmentation
We introduce the proposed double entropy data augmentation here. Shannon Entropy (Shannon 1948) quantifies the amount of information required to describe or encode a random variable’s possible states. In RL tasks, observations and trajectories are matrices in practice, but can also be normalized and viewed as discrete random variable. Consider a $m\times n$ matrix $A$ with each element $a_{ij}\geq 0,\,\forall i=1,...,m,\,j=1,...,n$, and the Shannon Entropy of matrix $A$ is defined by
$H(A):=-\sum_{k=1}^{mn}p_{k}\log(p_{k}),$ (6)
Denote state space as $\mathcal{S}\in\mathbb{R}^{m_{1}}$ and action space as $\mathcal{A}\in\mathbb{R}^{m_{2}}$, with $m_{1},\,m_{2}\geq 1$. Consider a trajectory $\tau$ of length $N$ as $\langle s_{0},a_{0},r_{0},s_{1},a_{1},r_{1},\dots,s_{N}\rangle$ with $a_{t}\in\mathcal{A},\ s_{t}\in\mathcal{S}$, for all $t=1,\dots,N$, and rewards are given following the environment’s original reward function $r_{t}=R(s_{t},a_{t},s^{\prime}_{t})$.
The set of all trajectories can be stacked into the matrix $\bar{\Gamma}=\begin{bmatrix}S,\,A,\,R\end{bmatrix}$ with $S\in\mathbb{R}^{N\times m_{1}},\,A\in\mathbb{R}^{N\times m_{2}}$ and $R\in\mathbb{R}^{N\times 1}$, and now we define the transform double entropy data augmentation $\sigma$ over matrix $\bar{\Gamma}$ as:
where $h^{n}\in\mathbb{R}$, equals to the entropy $H(s^{n})$, matrix $s^{n}$ is a submatrix of the stacked state $S$, obtained by equally dividing $S$ into $n$ parts along the state dimension. Here $n$ is a hyperparameter.
Appendix B Comparison of Data Augmentation Methods
(a) Seaquest
(b) Hero
(c) Montezuma’s Revenge
(d) Venture
Figure 7: Figures (a–d) demonstrate the best score (mean $\pm$ std over 3 seeds) in 1000 epochs of different data augmentation methods in 4 Atari environments. Note that Shannon refers to double entropy data augmentation. The legends of each sub figure are uniformly labeled in Figure(a)
Various types of data augmentation methods are used in the process of consistency regularization. Since the transformation double entropy data augmentation is thoroughly stated in previous section, we here list other transformation operations for data augmentation strategies involved in our experiment in Table 2 for completeness. Regarding to the heterogeneous SSRS variants which apply two weak and strong data augmentations in consistency regularization, the transformation details of the legends in Figures 3 in main text are listed below:
  • SSRS-S: SSRS-S applies random Gaussian noise with standard deviation $\sigma$ as the weak augmentation and double entropy data augmentation as strong augmentation.
  • SSRS-M: SSRS-M applies $\sigma$ random gaussian noise as the weak augmentation and smooth data augmentation as strong augmentation.
  • SSRS-C: SSRS-C applies $\sigma$ random gaussian noise as the weak augmentation and cutout data augmentation as strong augmentation.
Table 2: Data augmentation methods in the control group and SSRS variants. Weak augmentation is gaussian augmentation by default, and the rest augmentations are used as strong augmentations respectively.
Transformation Description Parameter Range
gaussian add $\sigma$ random gaussian noise to the observation $\sigma$ $default=0.1$
cutout randomly zero out $n$ columns of observation $n$ $default=16$
smooth smooth $n$ consecutive observations $n$ $default=3$
scale scale the observation by random scale factor $\lambda$ $\lambda$ (0.8,1.2)(0.8,1.2)( 0.8 , 1.2 )
translation horizontally translate the observation by factor $\lambda$ $\lambda$ (0,0.1)(0,0.1)( 0 , 0.1 )
flip horizontally mirror flip the observation
The experiments in the paper are conducted on a Windows platform with an Intel i5-13400F and an Nvidia RTX 5060. The dependencies and libraries required for each task can be found in the code appendix. Here we give the supplementary material and analysis regarding to the experiment section. Figure 7 presents the performance of SSRS when employing different data augmentation methods as SSL perturbation, including our proposed double entropy data augmentation (labeled as ”shannon” in the legend), along with conventional augmentation techniques such as flip, scale, and translate. In the first two games, Seaquest and Hero, double entropy data augmentation consistently maintains a leading position in terms of both convergence speed and final highest rewards, with the variance of the best reward remaining within a narrow range. In the more reward-sparse environment of MonteZumaRevenge, although the final best reward achieved by the translate augmentation (blue curve) method is higher, it exhibits significantly larger variance. In contrast, double entropy data augmentation (green curve) demonstrates much smaller variance (Figure 7(c)), indicating a more stable policy and a faster update rate. In Venture, double entropy data augmentation maintains stable variance throughout (Figure 7(d)), whereas other data augmentation methods struggle to avoid larger variance in such a sparse environment, where disturbance to the data exceeded the decision-making boundaries of the trajectory space. Moreover, double entropy data augmentation achieves a best score that was 10 percentage points higher than the best result among tested image-based data augmentation methods in this environment.
Appendix C Hyperparameters Settings
The hyperparameter setup of our experiments consists of the data augmentation parameters, the SSRS framework parameters and other general configurations. As for data augmentation parameters, we adopt consistent settings throughout the experiments. Due to different environments, the shapes of states and observations are different, so the parameter $n$ of double entropy data augmentation and cutout data augmentation, which is related to the state and observation dimension, are different in Atari and Robotic environments (see Table 3).
Value of $n$ Double-Ent Cutout
Atari 8 16
Robotic 4 3
In Table 4, we present the settings of hyperparameters in SSRS framework, which are based on preliminary hyperparameter tuning results. Note that dynamic setting is applied to the consistency coefficient $\alpha$ and the threshold of confidence $\lambda$, i.e., the value of these hyperparameters change during the training process, and the value shown in the Table 4 is the final value of the dynamic setting. In detail, the confidence threshold $\lambda$ grows along with the training process. In a training process with $T$ episodes in total, $\lambda$ is as followed at the $t$ episode:
$\lambda=0.6+0.3\times\left(1-e^{-\frac{t}{T}}\right)$
Similarly, in a training process with $T$ episodes in total, the consistency coefficient $\alpha$ is as followed at the $t$ episode: